A Complete Bibliography of Publications in Applied Mathematics and Computation: 1980–1984

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 December 2023 Version 1.06

Title word cross-reference

2 [Kle82]. 3 [CCSK82]. ^{N} [RUS^{+}84]. B [Rob82]. C^{3} [Tes80]. F (u(t), t, λ) = λ ∫_{0}^{1} k(t, y, u(y)) dy [KK83a]. n [CY80]. Q [Duc83].

-Analysis [Duc83]. -D [CCSK82, Kle82]. -dimensional [CY80].

12 [Ano83d, Ano84b]. 167 [Ano84b].

2- [Ano84b]. 205 [BB83b].

absolute [AK80]. abstract [Du83]. Accelerating [LS81]. acceleration [Joh83]. accumulations [DM84]. Acknowledgement [Ano82a]. Adaptive [Bra82, DSK82, AR82, Gno82, SB82, Van83]. adaptive-grid [Gno82]. advances [KB80]. airfoil [Sor82]. airfoils [Hal82]. algebra [KR81a]. Algebraic [Smi82, Stü83, Eis82, Sas83]. Algorithm [Kel80, AK80, Cro80, Gno82, IA80, KRTW83, KK83b, TZ83, Wat81, WI80]. Algorithms [Van83, GMS83, HK84]. AMG [Stü83]. Analysis [Var83, AHKK80, HG84, HEH81, Leo84, MH81, Wan81a, Wan81b, Yag82, Duc83]. Analytical [KR81a, Sas83]. Aperiodic [Wan81a]. applicable [Mus80]. Application [Kni82, Pap84, GD84, GT82, Kle82, Sas83, Smi84b, Ste82, Zie84]. Applications [MD81, SB82, GPAB83, McW82, Wat81]. Applied [Ano84b, Mus84, NSE81]. approach [Fie84, Lak81, ORSS83, Wei84].

Approximate [Kha81]. Approximation [GZ84]. approximations [AM81]. arbitrary [WZ82]. area [Wan80]. arising [LP81]. arithmetics [Mus80]. array [CCSK82]. aspects [KR81a, YR84]. Assessing [KK82]. atmospheric [Uen80]. Augmentor [Sor82]. Augmentor-Wing [Sor82]. Author [Ano83a]. Automated [MB82]. Automatic [Eis82, KS84a, KS84b, Rob82, KRT83]. autonomous [Sau83]. average [Kha81].

Banach [PV81]. based [MH80]. Basic [War82]. Bayes [HT80]. Bayesian [HW80]. behavior [dPL84]. between [WZ82]. Beyond [Ste84]. bifurcating [KK83b]. Bifurcation [dPL84, HEH81, HK84, KK83b, Sau83]. biharmonics [SSB82]. bilateral [KNS81]. Black [Den83]. block [Hil80]. board [Ano80, Ano82b, Ano83c, Ano81a, Ano81b, Ano83b, Ano84a]. bodies [WZ82]. body [Ste82]. boundaries [Lan81]. Boundary [Zam84, BKS84a, BKS84b, CCSK82, Col82, Dul82, Har82, Hau82, Joh82, LP81, LRV84, McW82, NSE81, Pen84, SV83, Tha82]. boundary-conforming [Dul82]. boundary-fitted [CCSK82, Col82, Hau82, Joh82, Tha82]. boundary-value [NSE81, Pen84]. bounded [Wan81b]. bounds [AK80, JTT80]. box [Den83]. branches [KK83b]. Brown [BK80].

C [Dul82]. C-type [Dul82]. C. [Ano83d, Ano84b]. Calculation [Lan81, Cau83, Jou82]. calculations [Abl82, Gno82]. calculus [KS84a]. capital [DM84, Puu84]. case [Wan81a]. cases [Wan81b]. catastrophe [Ste84]. Cauchy [AHKK80, Ioa83]. CDC [BB83b]. cell [MH81]. characterization [Smi84a]. chemical [HEH81]. chemotherapy [NHG83, OTA83]. choice [ABES84, Leo84, Smi84a]. Chow [Wat81]. Christoffel [ADHE82]. circular [CCSK82]. Classes [LS81]. Classic [Kel81, Kel80]. close [HKK^{+}80]. close-support [HKK^{+}80]. closed [WH84]. closed-form [WH84]. code [Jes83]. coefficients [BF83, TT84]. Collocation [Zam84, Ioa83]. communications [HKK^{+}80]. company [KPJ83]. Comparative [MH81, HK84]. Comparison [SV83, SV84, AM81, VV80]. comparisons [Stü83]. complementarity [Smi84b]. complex [HK84, WSL83, dPL84]. complex-valued [WSL83]. composite [Tho82a]. compressible [Yag82]. compromise [GPAB83]. Computation [Ano84b, Any82, Kni82]. Computational [YR84, Col82, Dul82, KR81a, KB80]. computations [WI80]. computer [GMS83, Har82, KT83, WI80, KT83]. condition [Smi84b]. conditions [BKS84a, BKS84b]. cones [Du83]. Confidence [JT83]. Conformal [Hal82, Har82, Ive82]. conforming [Dul82, Ste82]. conjugate [GMS83]. connected [Har82]. consistency [Var80]. constraints [KS81]. construction [Sas83]. Constructive [Pen84]. continuation [KK83b]. Continued [MB80]. Continued-fraction [MB80]. Continuous [RC81]. control [Bra82, KS84a, KS84b, ORSS83, YR84]. Convergence [Kel81, LS81, Joh83]. convexity [Kha81]. Coordinate [Bra82, CCSK82, Chr82, Col82, Eis82, Hau82, Joh82, KK82, Kni82, Mas82, McW82, RL82, Tha82, Tho82d, War82, Yag82]. coordinates [WZ82]. Cosine [Kel81, Kel80]. Courant [Pap84]. covariance [Sug83, Uen83]. criteria [SL83]. current [Mus84]. curve [MH81]. curves [KR81b, MH80]. curvilinear [ADHE82, KK82, Kni82, Ste82, Tha82, Tho82d]. Cyber [BB83b].

D [CCSK82, Kle82]. damped [HG84]. data [HK81, JT83, Van83]. Decell [KRTW83]. decentralized [Lak81]. decomposing [KK82]. degenerate [Uen83]. Degraded [HKK^{+}80]. delay [Du83, Sau83]. delivery [NHG83]. density [Kap80]. dependence [Kap80]. derivative [KRT83]. derivatives [KT83]. derived [OL80]. describing [PV81]. Design [Jes83, KS81, BB83a, Sug83]. Determination [KK83b]. Determining [Kru81]. deterministic [AM81, AS81]. development [HKWdZ83]. difference [DSK82, Ste82]. Differential [GZ84, AM81, AS81, Ado81, AR82, Cro80, Du83, HEH81, HK81, HK84, IM80, KST80, Nak82, Nie84, SL83, Sau83, SV84, Sor82, VV80, Vau80, War82, WSL83]. differentiation [HL82]. diffusion [GMS83]. digits [IRU83, RUS^{+}84]. dilemmas [Mus84]. dimensional [APYV83, BF83, Cau83, CY80, Dul82, Jam83, Jou82, MB82, MS83, OL80, Shi82, SSB82, Tha82, Tho82a, WZ82]. Dirac [APYV83]. Direction [KK83b]. Dirichlet [BKS84a]. discontinuous [BF83]. Discrete [CG83, AT80, HL82, YR84]. discrete-time [YR84]. discretized [BKS83]. discussion [Har82]. disease [MH81]. distribution [Duc83]. drug [NHG83]. drug-delivery [NHG83]. Dual [Kel81, Kel80]. dynamic [Abl82, Any82, VL83]. dynamical [KPJ83]. dynamics [Den84].

ecology [JTT80, JT80]. Editorial [Ano80, Ano82b, Ano83c, Ano81a, Ano81b, Ano83b, Ano84a]. effect [HT80]. efficient [WI80]. eigenvalue [Var80]. Element [Zam84, Sor82]. elementary [Ste84]. Elliptic [Tho82b, KB84, Kle82, Roa82, Sor82, Tha82, Tho82a, Van83]. Embedded [MS83]. embedding [Lor80]. energy [KS81]. Engineering [Wat81]. entrainment [OS80]. entropy [Fie84]. environmental [BB83a]. Equation [Kel81, APYV83, BKS83, BKS84a, BKS84b, KK83a, KST80, Kel80, OL80, Zie84]. Equations [GZ84, Zam84, AM81, AS81, Ado81, AHKK80, AR82, Cro80, Du83, HK84, Ioa83, Jam83, Jes83, KB84, Lor80, Nak82, Nie84, PV81, Pap84, SL83, Sas83, Sau83, SV84, Shi82, Sor82, TT84, VV80, Vau80, WH84]. Equidistant [Abl82]. equilibrium [Smi84b, Wei84]. equivalence [AHKK80, BK80]. Errata [Ano84b]. Error [HG84, Let80, Mas82]. estimates [HT80, JT83, KS80, Let80]. estimating [MH80]. estimation [BB83a, Uen83]. estimators [HW80, Sug83]. estuarine [Joh82]. Euler [Jam83, Jes83]. evaluated [KT83]. Evaluation [Duc83, HK84, KRT83]. evolving [ABES84]. exact [KT83]. Existence [Moo81, Nie84, VV80, Pen84]. expansions [MB80]. experience [Jou82]. experiences [Stü83]. experiments [OL80]. extension [BK80]. external [Ste82]. extreme [Smi84a].

failure [HT80]. family [KR81b]. Fast [Dul82, HKWdZ83]. field [Mus80]. finding [KST80]. Finite [Zam84, DSK82, Gno82, KRTW83, Ste82]. finite-volume [Gno82]. first [NSE81, Nie84, SV83]. fitted [CCSK82, Col82, Hau82, Joh82, McW82, Tha82]. Fitting [HK81]. flow [CCSK82, Jam83, Jou82, Ste82, Tro80, Wei84, Yag82]. flows [ADHE82, Cau83, Kle82, Kni82, MS83, Puu84]. Fokker [Zie84]. Forest [Cas83]. Foreword [Tho82c]. form [WH84]. fraction [MB80]. Fredholm [Lor80]. frequency [Wan81a, Wan81b]. frictions [Wei84]. function [HT80, KT83, Wan81a, Wan81b]. functions [HL82, IM80].

Galerkin [GZ84]. gas [Abl82]. Gaussian [Let80]. General [SL83, Tho82d, Kap80, Wei84]. generalised [Rob82]. generalizations [SB82]. Generalized [HW80, KRTW83, MB80, MD81, Smi84a]. generated [ADHE82, RUS^{+}84, WH84]. generating [SB82]. Generation [Col82, Tha82, Dul82, Eis82, GT82, Hal82, Ive82, Jou82, Kle82, Kni82, Nak82, Roa82, Rob82, Shi82, SSB82, Smi82, Sor82, Tho82a, Tho82b, War82, WZ82]. geometric [HK81]. geometric-mean [HK81]. geometries [Mus80, Tha82]. given [Sug83]. Global [Sau83]. gradient [Pen84, KS84b]. gradients [GMS83]. Grid [Sor82, Any82, BF83, Cau83, Gno82, GT82, Hal82, Ive82, Kle82, MB82, Nak82, Roa82, Shi82, SSB82, Smi82, Tho82b]. gridding [DSK82]. grids [ADHE82, AR82, Dul82, Ste82, Tho82a]. growth [Kap80, KPJ83, Puu84].

Haar [RJY80]. Hadamard [RJY80]. Harmonic [Wan81a, Wan81b]. harvesting [Cas83]. heat [DSK82, Lan81]. Helmholtz [OL80]. higher [KT83]. higher-order [KT83]. Hill [Kru81]. homogeneous [NSE81]. Hopf [HEH81]. housing [Leo84]. hydrodynamic [Pap84]. hydrodynamics [Joh82]. hypernumber [Mus80]. Hypernumbers [Ano83d, Ano84b, Mus80, Mus83].

Idealized [Any82]. identification [KS80]. II [BKS84b, Wan81b]. III [HKK^{+}80]. imbedding [Hil80, NSE81, Sug83]. implementation [Har82, Jes83]. implicit [AT82]. incomplete [Lak81]. Index [Ano82c, Ano83a, Var80]. induced [Mas82]. inequalities [AT80]. Infinite [LS81, CCSK82]. information [GRZ84, HW80, Lak81]. Initial [GZ84, KK83a, Uen83]. Initial-Value [GZ84, Uen83]. inputs [KS80]. integral [AHKK80, Ioa83, KK83a, Lor80]. Integrals [LS81]. Integration [Zam84, Wan80]. integrodifferential [TT84]. intelligence [HKK^{+}80]. intensity [HT80]. interference [Kle82]. internal [ADHE82, Kni82]. interpolation [HL82]. Interurban [DM84]. interval [HG83]. intervals [JT83]. Introduction [Den84, Leo84]. Invariant [Hil80, NSE81, Lor80]. inverses [KRTW83]. Inversion [Uen80]. inviscid [Joh83]. Isaacson [Pap84]. iterative [AM81, Du83, Moo81]. IV [KS84b].

Jacobian [Chr82, KK82].

kind [NSE81]. kinetic [OTA83].

labor [Puu84]. Laguerre [KNS81, Sum84]. land [Leo84]. latent [Lan81]. Layer [Zam84]. leading [IRU83, RUS^{+}84]. learning [Lak81]. least [AK80, HW80, KRT83, Sug83, Uen83]. least-squares [Sug83, Uen83]. limit [AS81]. Limiting [Vis80, Tro80]. line [Sug83]. Linear [GZ84, KR81a, KS81, Kha81, Pen84, Sug83, TZ83, Vau80]. loss [HT80].

MACSYMA [WH84].

MACSYMA-generated [WH84]. manageable [Chr82]. management [CW84, GRZ84]. Many [RC81, KT83]. Map [RC81]. mapping [Rob82, Wan80]. mappings [GT82, Har82]. Marching [Nak82, BKS83, BKS84a, BKS84b, KB84, Roa82]. Marquardt [GPAB83]. mass [DSK82]. Mathematics [Ano84b, Mus84]. matrices [Var83]. matrix [KST80, KK82, MD81, Sum84, WH84]. max [ORSS83, TZ83]. maximum [Fie84]. maximum-entropy [Fie84]. May [Ano84b]. mean [HK81]. mechanics [McW82]. mesh [Abl82, Jou82, MS83]. meshes [Bra82, Chr82, KK82, SB82]. Method [Zam84, AM81, HG83, Ioa83, Jam83, KK83a, KST80, KS84b, Kle82, LP81, LRV84, Lor80, MH80, PV81, Pap84, Sug83, Tha82]. methods [HG84, HKWdZ83, HL82, SB82]. methotrexate [NHG83]. min [ORSS83, TZ83]. min-max [ORSS83, TZ83]. mineral [GD84, LMS84]. minicomputer [MB82]. minimal [CG83]. minimize [KT83]. mixed [Moo81, BKS84b]. modal [HG84]. model [HK81, JTT80, JT80, Kap80, KPJ83, LMS84, Puu84]. Modeling [ABES84, CG83, Den84, Joh82, Tes80]. models [Leo84, Smi84a, TZ83, War82]. modification [GPAB83]. moment [JTT80]. monitoring [Cas83]. Monotone [Du83, LP81, LRV84, PV81, Moo81]. most [KR81b]. moving [Lan81]. MTBF [HT80]. Multi [BF83, Cau83]. Multi-grid [BF83, Cau83]. multielement [Hal82]. Multigrid [HKWdZ83, BB83b, Den83, GMS83, Jam83, Jes83, MS83, Stü83, Van83]. multiobjective [GD84]. multiple [Moo81]. multiply [Har82]. Musés [Ano83d, Ano84b].

natural [CW84]. natural-resource [CW84]. Navier [Gno82]. negative [Sas83]. Network [BB83a]. Neuromuscular [OS80]. Newton [HG83]. non [TZ83]. non-linear [TZ83]. Nonautonomous [IM80]. Nonlinear [KRT83, LS81, APYV83, Den84, HEH81, Hau82, HK81, JTT80, JT80, KK83a, LP81]. nonlinearities [Ado81]. Nonmonotonic [Kru81]. nonseparable [KB84]. nonsteady [Yag82]. nonsymmetric [Den83]. normed [Kha81]. note [Ulw81, Var80]. Numerical [AS81, HEH81, Joh82, OL80, Tho82a, WZ82, APYV83, Cro80, HK84, HL82, Ioa83, KB84, TT84].

observations [KS80]. on-line [Sug83]. one [APYV83]. one-dimensional [APYV83]. onto [Har82]. operations [HKK^{+}80]. operators [Ano83d, Ano84b, Moo81, Mus83]. optical [Uen80]. Optimal [HKK^{+}80, KS80, KS84a, KS84b, NHG83, YR84]. optimal-control [KS84a, KS84b]. optimization [KB80, VL83]. Order [GZ84, KT83, LRV84, Nie84, SV83]. ordinary [Cro80, HEH81, HK84, SL83]. Orthogonal [Chr82]. overrelaxation [GMS83].

Pages [Ano84b]. parabolic [Nak82, SV84]. Parameter [JTT80, KS80]. parameters [AK80]. Parametric [VL83]. partial [AR82, HEH81, Nak82, Sor82]. Patched [RL82]. patterns [ABES84, IRU83]. PEC [Cro80]. performance [Duc83]. Periodic [RC81, Wan81b, Dul82, LRV84, Nie84, Sau83]. Perron [KST80]. person [Lak81]. phase [Lan81]. phenomena [PV81]. physical [Any82, Mus84]. physically [Mus80]. Picard [AM81]. Planar [RC81]. Planck [Zie84]. point [KK83b, NSE81, Pen84]. Points [RC81, HK84, KK83b, Moo81]. Poisson [BB83b, BKS83, BKS84a, BKS84b, Shi82]. policies [Cas83]. polynomial [HL82]. polynomials [MB80]. population [DM84, HK81, Kap80]. positive [KST80]. potential [Cau83, Jou82]. preconditioned [GMS83]. Preface [MT83]. preliminary [Har82]. price [Smi84b]. prior [HW80]. probability [Smi84a]. Problem [GZ84, AHKK80, AK80, GMS83, IM80, KS84a, Lak81, Uen83, Vis80]. problems [AT82, BF83, Den83, DSK82, Hau82, KS84a, KS84b, LP81, LRV84, NSE81, ORSS83, Pen84, SV83, Smi84b, Yag82]. procedure [GD84]. processes [LP81, Vis80]. product [Ado81, WI80]. Profile [Kru81]. profiles [Uen80]. program [KT83]. proof [KRTW83]. propeller [Jou82]. properties [RUS^{+}84].

quadrature [Let80]. quality [KK82]. quantum [Mus80]. quasifixed [Moo81]. quasilinear [Tho82a]. questions [CW84].

radiance [Uen80]. radiation [OL80]. Rationale [GPAB83]. rationality [Tes80]. Rationalized [RJY80]. ratios [MB80]. reachability [CW84]. reactor [HEH81]. real [Wan81a, Wan81b]. reciprocal [Var83]. redundant [Ulw81]. Rees [Pap84]. refinement [MB82]. regimen [NHG83]. regional [Den84]. regions [Col82, Har82]. regression [AK80, HW80, TZ83]. regulators [KS81]. related [Wan80]. reliability [HT80, Ulw81]. representative [KR81b]. reservoir [GRZ84, Zie84]. resetting [OS80]. resonance [SV83, SV84]. resource [CW84]. resources [GD84, LMS84]. results [SV83, SV84, VV80]. Riccati [Sas83, WH84]. ridge [HW80]. Robinson [BK80]. root [KST80]. rotating [Jou82].

samples [Fie84]. sampling [CY80]. scheme [AT82]. Schwarz [ADHE82]. searching [IA80]. Seasonal [JT83]. Second [GZ84, LRV84, SV83]. Second-Order [GZ84]. segmented [Col82]. selection [GD84]. semidefinite [Sas83]. Semidirect [Roa82]. Semidirect/marching [Roa82]. Sensitivity [KS80, AHKK80, Var83]. Series [LS81]. sets [Kha81]. several [IM80]. sex [HK81]. shallow [Pap84]. shallow-water [Pap84]. shape [KR81b]. shapes [Har82, WZ82]. sickle [MH81]. similarity [Tro80]. Simplest [KS84a]. Simulation [JT80]. singular [Ioa83, TT84]. singularity [Kle82, NSE81]. Sobolev [VV80, Vau80]. solar [OL80, Uen80]. Solid [McW82]. Solution [ADHE82, Hau82, Jam83, AS81, AR82, BF83, CCSK82, Cro80, Ioa83, KB84, KS84a, KS84b, Sas83, Smi84b, Ste82, TT84, Tro80, Uen83]. Solutions [Kel81, Vau80, DSK82, Mus84, Nie84, Roa82, Sau83, WH84]. solve [Pap84]. solver [BB83b, Tha82]. solvers [HKWdZ83, Van83]. Solving [WSL83, AR82, GMS83]. Some [CW84, KB80, Mus84, AT80, Sau83, WH84]. space [OL80, PV81]. spaced [JT83]. spaces [Kha81]. spatial [ABES84, BB83a, Leo84, Puu84, Smi84b]. specifications [Sug83]. specified [Har82]. spectral [JT83]. spline [Rob82]. square [CCSK82]. squares [HW80, KRT83, Sug83, Uen83]. stability [DM84, IM80, Vis80]. starting [KK83b]. stochastic [AM81, AS81, Ado81, HW80, JTT80, JT80, Zie84]. stock [Wei84]. stock-flow [Wei84]. stocks [Puu84]. Stokes [Gno82]. storage [ORSS83]. String [IRU83, IA80]. strings [RUS^{+}84]. structural [DM84, HG84, RUS^{+}84]. structures [Van83]. study [APYV83, HT80, HK84, OTA83]. subpopulations [MH80]. Successive [GMS83, AM81]. summary [Mus80]. sums [Kha81]. support [HKK^{+}80]. surface [Wan80]. surfaces [CG83]. survival [MH80, MH81]. swirling [Tro80]. Sylvester [MD81]. Symmetric [BKS83, BKS84a, BKS84b, KB84]. symmetrical [Tes80]. system [AHKK80, Bra82, CCSK82, Duc83, HK81, Joh82, KS80, LRV84, LMS84]. systems [Any82, Col82, HG84, HEH81, Hau82, Hil80, IM80, Mas82, McW82, RL82, Tha82, Tho82a, Tho82d, Ulw81, VL83, WSL83, dPL84].

tactics [HKK^{+}80]. team [AHKK80, Lak81]. technique [BKS83, BKS84a, BKS84b, Du83, GD84, KB84, Yag82]. techniques [Kni82, Moo81]. term [Pen84]. Test [Yag82]. their [GT82, Mus80]. theorem [BK80]. theorems [CY80]. theory [AS81, HEH81, Mus80, Ste84, Zie84, AHKK80]. thickness [Uen80]. Three [Shi82, SSB82, BF83, Cau83, Dul82, Jou82, MB82, OL80, Tha82, Tho82a, WZ82].

Three-dimensional [Shi82, BF83, Cau83, Dul82, Jou82, MB82, OL80, Tha82, WZ82]. time [Ano83d, Ano84b, Mus83, YR84]. tip [Tha82]. topology [Rob82]. trade [Wei84]. transfer [DSK82, OL80]. Transfinite [GT82]. transform [KNS81, RJY80, Sum84]. transformation [ADHE82]. Transformations [LS81, Yag82]. transmitted [Uen80]. transonic [Cau83, Jam83, Kle82, MS83]. transport [LP81, PV81]. transportation [Leo84]. treatment [MS83]. trends [JT83]. tri [Sor82]. tri-element [Sor82]. tridiagonal [Hil80]. tubes [CCSK82]. Two [LS81, HK81, Jam83, Lak81, MS83, NSE81, Pen84]. two-dimensional [MS83]. two-point [NSE81, Pen84]. two-sex [HK81]. type [Dul82, Ioa83, VV80].

unconstrained [KB80]. undersized [Fie84]. unequally [JT83]. uniqueness [Pen84, SL83]. upon [MH80]. urban [Den84]. use [AR82, Leo84]. using [CCSK82, Col82, Hau82, HL82, KT83, Kle82, Nak82, Shi82, SSB82, Tha82]. utilization [LMS84].

Validation [Lor80]. Value [GZ84, AK80, GRZ84, KK83a, LP81, LRV84, NSE81, Pen84, SV83, Smi84a, Uen83]. valued [WSL83]. variable [TT84]. variables [BB83a, KT83]. variational [AHKK80, SB82]. variations [KS84a]. vector [GMS83]. Vectorized [BB83b, Gno82]. via [KRT83, Sug83]. viscous [ADHE82, Joh83]. viscously [HG84]. volume [Gno82]. volumes [Ano83a]. vortex [Tro80].

water [AT82, GD84, Hau82, Pap84]. wave [AT82, Hau82]. wing [Tha82, Sor82]. wing-tip [Tha82]. wing/wing [Tha82]. wing/wing-tip [Tha82].

Yorke [Wat81].

References

Allen:1984:MES

[ABES84] P. M. Allen, F. Boon, G. Engelen, and M. Sanglier. Modeling evolving spatial choice patterns. Applied Mathematics and Computation, 14(1):97–129, January 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                                   sciencedirect.com/science/
                                                   article/pii/0096300384900481.

Ablow:1982:EMG

[Abl82] C. M. Ablow. Equidistant mesh for gas dynamic calculations. Applied Mathematics and Computation, 10–11:859–863, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/0096300382902260.

Anderson:1982:SVI

[ADHE82] O. L. Anderson, R. T. Davis, G. B. Hankins, and D. E. Edwards. Solution of viscous internal flows on curvilinear grids generated by the Schwarz— Christoffel transformation. Applied Mathematics and Computation, 10–11:507–524, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/0096300382902053.

Adomian:1981:PNS

[Ado81] G. Adomian. On product nonlinearities in stochastic differential equations. Applied Mathematics and Computation, 8 (1):35–49, January 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900333.

Akbari:1980:ETT

[AHKK80] Alireza Akbari, James Hess, Harriet Kagiwada, and Robert Kalaba. The equivalence of team theory’s integral equations and a Cauchy system: sensitivity analysis of a variational problem. Applied Mathematics and Computation, 6(1):21–36, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900132.

Armstrong:1980:ALA

[AK80] Ronald D. Armstrong and Mabel Tam Kung. An algorithm for a least absolute value regression problem with bounds on the parameters. Applied Mathematics and Computation, 7(3): 267–279, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900491.

Adomian:1981:CIM

[AM81] G. Adomian and K. Malakian. A comparison of the iterative method and Picard’s successive approximations for deterministic and stochastic differential equations. Applied Mathematics and Computation, 8 (3):187–204, April 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381900060.

Anonymous:1980:EB

[Ano80] Anonymous. Editorial board. Applied Mathematics and Computation, 7(1):??, July 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300380900296.

Anonymous:1981:EBa

[Ano81a] Anonymous. Editorial Board. Applied Mathematics and Computation, 8(1):??, January 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381900308.

Anonymous:1981:EBb

[Ano81b] Anonymous. Editorial Board. Applied Mathematics and Computation, 9(1):??, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/009630038190120X.

Anonymous:1982:A

[Ano82a] Anonymous. Acknowledgement.

Applied Mathematics and Computation, 10–11:xvii, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901837.

Anonymous:1982:EB

[Ano82b] Anonymous. Editorial board.

Applied Mathematics and Computation, 10–11:iii–iv, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901825.

Anonymous:1982:I

[Ano82c] Anonymous.  Index.  Ap-

plied Mathematics and Computation, 10–11:895–909, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902296.

Anonymous:1983:AIV

[Ano83a] Anonymous. Author index to volumes 12 and 13, 1983. Applied Mathematics and Computation, 13(3–4):471–473, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900255.

Anonymous:1983:EBa

[Ano83b] Anonymous. Editorial Board. Applied Mathematics and Computation, 12(1):i–ii, February 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900383.

Anonymous:1983:EBb

[Ano83c] Anonymous. Editorial board. Applied Mathematics and Computation, 13(1–2):iii–iv, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900267.

Anonymous:1983:HTO

[Ano83d] Anonymous. Hypernumbers and time operators: C. Musés, 12:139–168 (1983). Applied Mathematics and Computation, 12(4):395, July 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900498.

Anonymous:1984:EB

[Ano84a] Anonymous. Editorial Board. Applied Mathematics and Computation, 15(1):ii–iii, July 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900493.

Anonymous:1984:ECM

[Ano84b] Anonymous. Errata: C. Musès, “Hypernumbers and time operators”, Applied Mathematics and Computation, 12(2-?3), May 1983, Pages 139–167. Applied Mathematics and Computation, 14(2):219, February 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900390.

See [Mus83].

Anyiwo:1982:IDG

[Any82]  Joshua C. Anyiwo.  Ideal-

ized dynamic grid computation of physical systems. Applied Mathematics and Computation, 10–11:837–858, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902259.

Alvarez:1983:NSN

[APYV83] A. Alvarez, Kuo Pen-Yu, and Luis Vasquez. The numerical study of a nonlinear onedimensional Dirac equation. Applied Mathematics and Computation, 13(1–2):1–15, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900279.

Anderson:1982:USA

[AR82] Dale A. Anderson and M. M.

Rai. The use of solution adaptive grids in solving partial differential equations. Applied Mathematics and Computation, 10–11:317–338, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901965.

Adomian:1981:NSD

[AS81] G. Adomian and D. Sarafyan. Numerical solution of differential equations in the deterministic limit of stochastic theory. Applied Mathematics and Computation, 8(2):111–119, March 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900023.

Agarwal:1980:SND

[AT80] Ravi P. Agarwal and E. Thandapani. On some new discrete inequalities. Applied Mathematics and Computation, 7(3): 205–224, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900442.

Aston:1982:ISW

[AT82] Martha B. Aston and J. W. Thomas. An implicit scheme for water wave problems. Applied Mathematics and Computation, 10–11:809–818, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902235.

Bardossy:1983:NDS

[BB83a] A. Bárdossy and I. Bogárdi.

Network design for the spatial estimation of environmental variables. Applied Mathematics and Computation, 12 (4):339–365, July 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900462.

Barkai:1983:VMP

[BB83b] D. Barkai and A. Brandt. Vectorized multigrid Poisson solver for the CDC Cyber 205. Applied Mathematics and Computation, 13(3–4):215–227, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900139.

Behie:1983:MGS

[BF83] A. Behie and P. A. Forsyth, Jr. Multi-grid solution of threedimensional problems with discontinuous coefficients. Applied Mathematics and Computation, 13(3–4):229–240, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900140.

Brown:1980:EBR

[BK80] Donald J. Brown and M. Ali Khan. An extension of the Brown–Robinson equivalence theorem. Applied Mathematics and Computation, 6(2): 167–175, March 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300380900405.

Bharadwaj:1983:SMT

[BKS83] K. K. Bharadwaj, M. K. Kadalbajoo, and R. Sankar. Symmetric marching technique for the discretized Poisson equation. Applied Mathematics and Computation, 12(2–3):187–198, May 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900061.

Bharadwaj:1984:SMTa

[BKS84a] Kamal K. Bharadwaj, Mohan K. Kadalbajoo, and R. Sankar. Symmetric marching technique for the Poisson equation. I. Dirichlet boundary conditions. Applied Mathematics and Computation, 15(2):137–149, August 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300384900134.

Bharadwaj:1984:SMTb

[BKS84b] Kamal K. Bharadwaj, Mohan K. Kadalbajoo, and R. Sankar.

Symmetric marching technique for the Poisson equation. II. Mixed boundary conditions. Applied Mathematics and Computation, 15(2):151–163, August 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300384900146.

Brackbill:1982:CSC

[Bra82] J. U. Brackbill. Coordinate system control: Adaptive meshes. Applied Mathematics and Computation, 10–11:277–294, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901941.

Casti:1983:FMH

[Cas83] J. Casti. Forest monitoring and harvesting policies. Applied Mathematics and Computation, 12(1):19–48, February 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900401.

Caughey:1983:MGC

[Cau83] D. A. Caughey. Multi-grid calculation of three-dimensional transonic potential flows. Applied Mathematics and Computation, 13(3–4):241–260, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900152.

Chen:1982:DSF

[CCSK82] B. C-J. Chen, T. H. Chien, W. T. Sha, and J. H. Kim. 3-D solution of flow in an infinite square array of circular tubes by using boundaryfitted coordinate system. Applied Mathematics and Computation, 10–11:619–632, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382902119.

Coppin:1983:DMM

[CG83] C. Coppin and D. Greenspan. Discrete modeling of minimal surfaces. Applied Mathematics and Computation, 13(1– 2):17–35, August 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900280.

Christov:1982:OCM

[Chr82] C. I. Christov. Orthogonal coordinate meshes with manageable Jacobian. Applied Mathematics and Computation, 10–11:885–894, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382902284.

Coleman:1982:GBF

[Col82] Roderick M. Coleman. Generation of boundary-fitted coordinate systems using segmented computational regions. Applied Mathematics and Computation, 10–11:633–651, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902120.

Crown:1980:NPA

[Cro80] J. C. Crown. A new PEC algorithm for the numerical solution of ordinary differential equations. Applied Mathematics and Computation, 6(3):189–209, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300380900508.

Casti:1984:SQR

[CW84] J. Casti and E. F. Wood.

Some questions of reachability in natural-resource management. Applied Mathematics and Computation, 15(3): 185–207, October 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900018.

Chen:1980:DST

[CY80] K. H. Chen and C. C. Yang.

On n-dimensional sampling theorems. Applied Mathematics and Computation, 7(3):

247–257, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900478.

Dendy:1983:BBM

[Den83] J. E. Dendy, Jr. Black box multigrid for nonsymmetric problems. Applied Mathematics and Computation, 13(3–4):261–283, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900164.

Dendrinos:1984:IND

[Den84] D. S. Dendrinos. Introduction — nonlinear dynamics in urban and regional modeling. Applied Mathematics and Computation, 14(1):1–2, January 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900407.

Dendrinos:1984:IPC

[DM84] Dimitrios S. Dendrinos and Henry Mullally. Interurban population and capital accumulations and structural stability. Applied Mathematics and Computation, 14(1):11–24, January 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300384900420.

dePalma:1984:BBC

[dPL84] André de Palma and Claude Lefèvre. Bifurcation and behavior of complex systems. Applied Mathematics and Computation, 14(1):77–95, January 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038490047X.

Dwyer:1982:AGF

[DSK82] Harry A. Dwyer, Mitchell D.

Smooke, and Robert J. Kee. Adaptive gridding for finite difference solutions to heat and mass transfer problems. Applied Mathematics and Computation, 10–11:339–356, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901977.

Du:1983:MIT

[Du83] S. W. Du. Monotone iterative technique for delay differential equations in abstract cones. Applied Mathematics and Computation, 12(2–3):213–219, May 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900085.

Duckstein:1983:EPD

[Duc83] Lucien Duckstein. Evaluation of the performance of a distribution system by Q-Analysis. Applied Mathematics and Computation, 13(1–2):173–184, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038390036X.

Dulikravich:1982:FGT

[Dul82] Djordje S. Dulikravich. Fast generation of three-dimensional computational boundary-conforming periodic grids of C-type. Applied Mathematics and Computation, 10–11:563–584, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902089.

Eiseman:1982:AAC

[Eis82] Peter R. Eiseman. Automatic algebraic coordinate generation. Applied Mathematics and Computation, 10–11:447–463, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902028.

Fiebig:1984:MEA

[Fie84]  Denzil G. Fiebig.  On the

maximum-entropy approach to undersized samples. Applied Mathematics and Computation, 14(3):301–312, April 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900274.

Gershon:1984:PSM

[GD84] Mark Gershon and Lucien Duckstein. A procedure for selection of a multiobjective technique with application to water and mineral resources. Applied Mathematics and Computation, 14(3):245–271, April 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900249.

Gary:1983:SOM

[GMS83] John Gary, Steve McCormick, and Roland Sweet. Successive overrelaxation, multigrid, and preconditioned conjugate gradients algorithms for solving a diffusion problem on a vector computer. Applied Mathematics and Computation, 13(3–4): 285–309, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900176.

Gnoffo:1982:VFV

[Gno82] Peter A. Gnoffo. A vectorized, finite-volume, adaptivegrid algorithm for Navier– Stokes calculations. Applied Mathematics and Computation, 10–11:819–835, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902247.

Garcia-Pena:1983:MMC

[GPAB83] J. Garcia-Peña, S. P. Azen, and R. N. Bergman. On a modification of Marquardt’s compromise: Rationale and applications. Applied Mathematics and Computation, 12(1):1–17, February 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900395.

Guariso:1984:VIR

[GRZ84] G. Guariso, S. Rinaldi, and P. Zielinski. The value of information in reservoir management. Applied Mathematics and Computation, 15(2):165–184, August 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300384900158.

Gordon:1982:TMT

[GT82] William J. Gordon and Linda C. Thiel. Transfinite mappings and their application to grid generation. Applied Mathematics and Computation, 10–11:171–233, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300382901916.

Gregory:1984:GAI

[GZ84] John Gregory and Marvin Zeman. A Galerkin approximation for the initial-value problem for linear second-order differential equations. Applied Mathematics and Computation, 15 (2):93–108, August 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900109.

Halsey:1982:CGG

[Hal82] Douglas Halsey. Conformal grid generation for multielement airfoils. Applied Mathematics and Computation, 10–11:585–600, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300382902090.

Harrington:1982:CMM

[Har82] Andrew Harrington. Conformal mappings onto multiply connected regions with specified boundary shapes: a preliminary discussion of computer implementation. Applied Mathematics and Computation, 10–11:601–618, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902107.

Haussling:1982:SNW

[Hau82]  Henry J. Haussling.  Solu-

tion of nonlinear water wave problems using boundary-fitted coordinate systems. Applied Mathematics and Computation, 10–11:385–407, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382901990.

Hassard:1981:NHB

[HEH81] B. Hassard and I. M. El-Henawy. Numerical Hopf bifurcation analysis in nonlinear ordinary and partial differential systems from chemical reactor theory. Applied Mathematics and Computation, 9(2):75–92, September 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300381900084.

Hansen:1983:INM

[HG83] E. R. Hansen and R. I. Greenberg. An interval Newton method. Applied Mathematics and Computation, 12(2– 3):89–98, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900012.

Hanson:1984:EAM

[HG84] P. M. Hanson and I. Gladwell. Error analysis of modal methods for viscously damped structural systems. Applied Mathematics and Computation, 14(2): 167–195, February 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900365.

Hillion:1980:IIB

[Hil80] P. Hillion. Invariant imbedding for block tridiagonal systems. Applied Mathematics and Computation, 6(2):95–122, March 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300380900363.

Haynie:1981:FND

[HK81] Susan P. Haynie and Robert E.

Kalaba. Fitting a nonlinear differential system to population data: the two-sex geometricmean model. Applied Mathematics and Computation, 9 (1):1–26, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381901211.

Holodniok:1984:NAE

[HK84] M. Holodniok and M. Kubiček.

New algorithms for the evaluation of complex bifurcation points in ordinary differential equations. A comparative numerical study. Applied Mathematics and Computation, 15(3): 261–274, October 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900055.

Hess:1980:OTC

[HKK^{+}80] J. Hess, Harriet Kagiwada, R. Kalaba, K. Spingarn, and C. Tsokos. Optimal tactics for close-support operations — III. Degraded intelligence and communications. Applied Mathematics and Computation, 6(3): 217–228, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900521.

Hemker:1983:MMD

[HKWdZ83] P. W. Hemker, R. Kettler, P. Wesseling, and P. M. de Zeeuw. Multigrid methods: development of fast solvers. Applied Mathematics and Computation, 13(3–4):311–326, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900188.

Hyman:1982:NDD

[HL82] James M. Hyman and Bernard Larrouturou. The numerical differentiation of discrete functions using polynomial interpolation methods. Applied Mathematics and Computation, 10–11:487–506, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902041.

Higgins:1980:SEL

[HT80] J. J. Higgins and C. P. Tsokos. A study of the effect of the loss function on Bayes estimates of failure intensity, MTBF, and reliability. Applied Mathematics and Computation, 6(2): 145–166, March 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900399.

Haitovsky:1980:GRR

[HW80] Yoel Haitovsky and Yohanan Wax. Generalized ridge regression, least squares with stochastic prior information, and Bayesian estimators. Applied Mathematics and Computation, 7(2):125–154, September 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900028.

Iyengar:1980:SSA

[IA80] S. Sitharama Iyengar and Vincent Alia. A string searching algorithm. Applied Mathematics and Computation, 6(2): 123–131, March 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900375.

Ianiro:1980:NDS

[IM80] N. Ianiro and C. Maffei. Nonautonomous differential systems: several functions in the stability problem. Applied Mathematics and Computation, 7 (1):71–80, July 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038090034X.

Ioakimidis:1983:NSC

[Ioa83] N. I. Ioakimidis. On the numerical solution of Cauchy type singular integral equations by the collocation method. Applied Mathematics and Computation, 12(1):49–60, February 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900413.

Iyengar:1983:SPL

[IRU83] S. Sitharama Iyengar, A. K. Rajagopal, and V. R. R. Uppuluri. String patterns of leading digits. Applied Mathematics and Computation, 12(4):321–337, July 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900450.

Ives:1982:CGG

[Ive82]  David C. Ives.  Confor-

mal grid generation. Applied Mathematics and Computation, 10–11:107–135, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382901898.

Jameson:1983:SEE

[Jam83] Antony Jameson. Solution of the Euler equations for two dimensional transonic flow by a multigrid method. Applied Mathematics and Computation, 13(3– 4):327–355, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038390019X.

Jespersen:1983:DIM

[Jes83] Dennis C. Jespersen. Design and implementation of a multigrid code for the Euler equations. Applied Mathematics and Computation, 13(3–4):357–374, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900206.

Johnson:1982:NME

[Joh82] Billy H. Johnson. Numerical modeling of estuarine hydrodynamics on a boundaryfitted coordinate system. Applied Mathematics and Computation, 10–11:409–436, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902004.

Johnson:1983:CAV

[Joh83]  Gary M. Johnson.  conver-

gence acceleration of viscous and inviscid. Applied Mathematics and Computation, 13(3– 4):375–398, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900218.

Jou:1982:EMG

[Jou82] Wen-Huei Jou. An experience in mesh generation for threedimensional calculation of potential flow around a rotating propeller. Applied Mathematics and Computation, 10–11:547–561, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300382902077.

Jernigan:1980:SNS

[JT80] Robert W. Jernigan and Chris P. Tsokos. Simulation of a nonlinear stochastic ecology model. Applied Mathematics and Computation, 7(1):9–25, July 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900314.

Jernigan:1983:STU

[JT83] Robert W. Jernigan and John C. Turner. Seasonal trends in unequally spaced data: Confidence intervals for spectral estimates. Applied Mathematics and Computation, 12(2–3):221–236, May 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900097.

Jernigan:1980:PMB

[JTT80] R. W. Jernigan, J. C. Turner, and C. P. Tsokos. Parameter and moment bounds for a nonlinear stochastic ecology model. Applied Mathematics and Computation, 7(1):27–53, July 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900326.

Kapur:1980:GPG

[Kap80] J. N. Kapur. A general population growth model with density dependence. Applied Mathematics and Computation, 6 (1):37–50, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900144.

Kowalik:1980:SCA

[KB80] J. S. Kowalik and W. R. Boland.

Some computational advances in unconstrained optimization. Applied Mathematics and Computation, 7(1):55–69, July 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900338.

Kadalbajoo:1984:SMT

[KB84] Mohan K. Kadalbajoo and K. K.

Bharadwaj. Symmetric marching technique for the numerical solution of nonseparable elliptic equations. Applied Mathematics and Computation, 14(2): 197–206, February 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900377.

Kelman:1980:ACD

[Kel80]  Robert B. Kelman.  Algo-

rithm for the classic dual cosine equation. Applied Mathematics and Computation, 6 (1):7–19, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900120.

Kelman:1981:CSC

[Kel81] Robert B. Kelman. Convergence of solutions to the classic dual cosine equation. Applied Mathematics and Computation, 9(3): 217–225, October 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900291.

Khan:1981:ACA

[Kha81] M. Ali Khan. Approximate convexity of average sums of sets in normed linear spaces. Applied Mathematics and Computation, 9(1):27–34, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381901223.

Kerlick:1982:AQC

[KK82] G. David Kerlick and Goetz H.

Klopfer. Assessing the quality of curvilinear coordinate meshes by decomposing the Jacobian matrix. Applied Mathematics and Computation, 10–11:787–807, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300382902223.

Kagiwada:1983:IVM

[KK83a] Harriet H. Kagiwada and Robert E.

Kalaba. An initial value method for the nonlinear integral equation F (u(t), t, λ) = λ ∫_{0}^{1} k(t, y, u(y)) dy. Applied Mathematics and Computation, 13(1–2):117–124, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900334.

Kubicek:1983:DBB

[KK83b] Milan Kubíček and Alois Klíč.

Direction of branches bifurcating at a bifurcation point. Determination of starting points for a continuation algorithm. Applied Mathematics and Computation, 13(1–2):125–142, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300383900346.

Klevenhusen:1982:DEG

[Kle82] Karl D. Klevenhusen. 2-D elliptic grid generation using a singularity method and its application to transonic interference flows. Applied Mathematics and Computation, 10–11:739–760, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382902193.

Knight:1982:ACC

[Kni82] Doyle D. Knight. Application of curvilinear coordinate generation techniques to the computation of internal flows. Applied Mathematics and Computation, 10–11:357–384, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382901989.

Keilson:1981:BLT

[KNS81] J. Keilson, W. Nunn, and U. Sumita. The bilateral Laguerre transform. Applied Mathematics and Computation, 8(2): 137–174, March 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381900047.

Karmeshu:1983:DMC

[KPJ83] Karmeshu, R. K. Pathria, and V. P. Jain. A dynamical model of company growth. Applied Mathematics and Computation, 12(1):61–75, February 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900425.

Kalaba:1981:AAC

[KR81a] Robert Kalaba and Nima Rasakhoo. Analytical aspects of computational linear algebra. Applied Mathematics and Computation, 8(2):83–109, March 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300381900011.

Kalaba:1981:MRS

[KR81b] Robert Kalaba and Nima Rasakhoo. The most representative shape of a family of curves. Applied Mathematics and Computation, 8(3):205–236, April 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300381900072.

Kalaba:1983:NLS

[KRT83] Robert Kalaba, Nima Rasakhoo, and Asher Tishler. Nonlinear least squares via automatic derivative evaluation. Applied Mathematics and Computation, 12(2–3):119–137, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900036.

Kalaba:1983:NPD

[KRTW83] Robert Kalaba, Nima Rasakhoo, Asher Tishler, and Jia Song Wang. A new proof for Decell’s finite algorithm for generalized inverses. Applied Mathematics and Computation, 12 (2–3):199–211, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900073.

Krueger:1981:DPN

[Kru81]  Robert J. Krueger.  Deter-

mining the profile of a nonmonotonic hill. Applied Mathematics and Computation, 8 (1):1–16, January 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038190031X.

Kalaba:1980:SPE

[KS80] R. E. Kalaba and K. Spingarn. Sensitivity of parameter estimates to observations, system identification, and optimal inputs. Applied Mathematics and Computation, 7(3): 225–235, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900454.

Kalaba:1981:DLR

[KS81] R. Kalaba and K. Spingarn. Design of linear regulators with energy constraints. Applied Mathematics and Computation, 8 (3):175–185, April 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900059.

Kalaba:1984:ASOa

[KS84a] R. Kalaba and K. Spingarn.

Automatic solution of optimalcontrol problems. I. Simplest problem in the calculus of variations. Applied Mathematics and Computation, 14(2): 131–148, February 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900341.

Kalaba:1984:ASOb

[KS84b] R. Kalaba and K. Spingarn.

Automatic solution of optimalcontrol problems. IV. Gradient method. Applied Mathematics and Computation, 14 (3):289–300, April 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900262.

Kalaba:1980:NDE

[KST80] R. Kalaba, K. Spingarn, and L. Tesfatsion. A new differential equation method for finding the Perron root of a positive matrix. Applied Mathematics and Computation, 7(3):187–193, October 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300380900429.

Kalaba:1983:CPM

[KT83] R. Kalaba and A. Tishler. A Computer program to minimize a function with many variables using computer evaluated exact higher-order derivatives. Applied Mathematics and Computation, 13(1–2):143–172, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300383900358.

Lakshmivarahan:1981:LAT

[Lak81] S. Lakshmivarahan. A learning approach to the two person decentralized team problem with incomplete information. Applied Mathematics and Computation, 8(1):79–82, January 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381900357.

Langerholc:1981:CMP

[Lan81] J. Langerholc. Calculation of moving phase boundaries with latent heat. Applied Mathematics and Computation, 9 (1):43–51, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381901247.

Leonardi:1984:ISC

[Leo84] Giorgio Leonardi. Introduction — spatial choice models in housing, transportation, and land use analysis. Applied Mathematics and Computation, 14 (1):33–34, January 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900444.

Lether:1980:EEG

[Let80] Frank G. Lether. Error estimates for Gaussian quadrature. Applied Mathematics and Computation, 7(3):237–246, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900466.

Lengyel:1984:SMU

[LMS84] Laszlo Lengyel, Sandor Molnar, and Ferenc Szidarovszky. A system model for the utilization of mineral resources. Applied Mathematics and Computation, 14(3):273–287, April 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300384900250.

Lord:1980:VIE

[Lor80] M. E. Lord. Validation of an invariant embedding method for Fredholm integral equations. Applied Mathematics and Computation, 7(1):1–7, July 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300380900302.

Lakshmikantham:1981:MMN

[LP81] V. Lakshmikantham and B. G. Pachpatte. Monotone method for nonlinear boundary value problems arising in transport processes. Applied Mathematics and Computation, 9 (1):35–41, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381901235.

Lakshmikantham:1984:MMS

[LRV84] V. Lakshmikantham, V. Sree Hari Rao, and A. S. Vatsala. Monotone method for a system of second order periodic boundary value problems. Applied Mathematics and Computation, 15(1):71–83, July 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900547.

Levin:1981:TNC

[LS81] David Levin and Avram Sidi.

Two new classes of nonlinear transformations for accelerating the convergence of infinite integrals and series. Applied Mathematics and Computation, 9(3): 175–215, October 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038190028X.

Mastin:1982:EIC

[Mas82] C. Wayne Mastin. Error induced by coordinate systems. Applied Mathematics and Computation, 10–11:31–40, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901862.

Maroulas:1980:CFE

[MB80] John Maroulas and Stephen Barnett. Continued-fraction expansions for ratios of generalized polynomials. Applied Mathematics and Computation, 6(3): 229–249, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900533.

Manhardt:1982:ATD

[MB82] P. D. Manhardt and A. J. Baker.

Automated three-dimensional grid refinement on a minicomputer. Applied Mathematics and Computation, 10–11:437–446, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300382902016.

McWhorter:1982:SMA

[McW82] John C. McWhorter. Solid mechanics applications of boundary fitted coordinate systems. Applied Mathematics and Computation, 10–11:253–276, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038290193X.

Maroulas:1981:AGS

[MD81] J. Maroulas and D. Dascalopoulos. Applications of the generalized Sylvester matrix. Applied Mathematics and Computation, 8(2):121–135, March 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900035.

Murthy:1980:NME

[MH80] Vrudhula K. Murthy and L. Julian Haywood. New method for estimating survival curves based upon subpopulations. Applied Mathematics and Computation, 7(3):259–266, October 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038090048X.

Murthy:1981:CSC

[MH81] Vrudhula K. Murthy and L. Julian Haywood. Comparative survival curve analysis in sickle cell disease. Applied Mathematics and Computation, 9(2): 143–152, September 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900126.

Moore:1981:EMQ

[Moo81] Jane Moore. Existence of multiple quasifixed points of mixed monotone operators by iterative techniques. Applied Mathematics and Computation, 9(2): 135–141, September 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900114.

McCarthy:1983:EMM

[MS83] D. R. McCarthy and R. C. Swanson. Embedded mesh multigrid treatment of two-dimensional transonic flows. Applied Mathematics and Computation, 13(3– 4):399–417, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038390022X.

McCormick:1983:P

[MT83] Steve McCormick and Ulrich Trottenberg. Preface. Applied Mathematics and Computation, 13(3–4):213, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900127.

Muses:1980:HQF

[Mus80] C. Musès. Hypernumbers and quantum field theory with a summary of physically applicable hypernumber arithmetics and their geometries. Applied Mathematics and Computation, 6(1):63–94, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900168.

Muses:1983:HTO

[Mus83] C. Musès. Hypernumbers and time operators. Applied Mathematics and Computation, 12 (2–3):139–167, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900048.

See errata [Ano84b].

Muses:1984:SCD

[Mus84] C. Musès. Some current dilemmas in applied physical mathematics with some solutions. Applied Mathematics and Computation, 14(2):207–217, February 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300384900389.

Nakamura:1982:MGG

[Nak82] S. Nakamura. Marching grid generation using parabolic partial differential equations. Applied Mathematics and Computation, 10–11:775–786, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902211.

Nathanson:1983:TOD

[NHG83] M. H. Nathanson, R. S. Hillman, and C. Georgakis. Towards an optimal drug-delivery regimen for methotrexate chemotherapy. Applied Mathematics and Computation, 12(2–3):99–117, May 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900024.

Nieto:1984:EPS

[Nie84] Juan J. Nieto. Existence of periodic solutions for first order differential equations. Applied Mathematics and Computation, 15(3):221–232, October 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300384900031.

Nelson:1981:IIA

[NSE81] Paul Nelson, Seok Sagong, and Ira T. Elder. Invariant imbedding applied to homogeneous two-point boundary-value problems with a singularity of the first kind. Applied Mathematics and Computation, 9(2):93– 110, September 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900096.

Ou:1980:NEH

[OL80] Szu-Cheng S. Ou and Kuo-Nan Liou. Numerical experiments on the Helmholtz equation derived from the solar radiation transfer equation in threedimensional space. Applied Mathematics and Computation, 7(2):155–175, September 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038090003X.

Orlovski:1983:MMA

[ORSS83] S. Orlovski, S. Rinaldi, and R. Soncini-Sessa. A minmax approach to storage control problems. Applied Mathematics and Computation, 12 (2–3):237–254, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900103.

Oguztoreli:1980:NRE

[OS80] M. N. Oğuztöreli and R. B. Stein. Neuromuscular resetting and entrainment. Applied Mathematics and Computation, 7(2): 93–124, September 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900016.

Oguztoreli:1983:KSC

[OTA83] M. N. Oğuztöreli, C. P. Tsokos, and J. Akabutu. A kinetic study of chemotherapy. Applied Mathematics and Computation, 12 (2–3):255–300, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900115.

Papa:1984:ACI

[Pap84] L. Papa. Application of the Courant–Isaacson–Rees method to solve the shallow-water hydrodynamic equations. Applied Mathematics and Computation, 15(1):85–92, July 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900559.

Pennline:1984:CEU

[Pen84] James A. Pennline. Constructive existence and uniqueness for twopoint boundary-value problems with a linear gradient term. Applied Mathematics and Computation, 15(3):233–260, October 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300384900043.

Puu:1984:MSF

[Puu84] Tönu Puu. A model of spatial flows and growth of capital and labor stocks. Applied Mathematics and Computation, 14(1):3–9, January 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900419.

Pachpatte:1981:MME

[PV81] B. G. Pachpatte and A. S. Vatsala. Monotone method for equations describing transport phenomena in a Banach space. Applied Mathematics and Computation, 9(3):165–174, October 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300381900278.

Rogers:1981:CPM

[RC81] Thomas D. Rogers and Bruce L. Clarke. A continuous planar map with many periodic points. Applied Mathematics and Computation, 8(1):17–33, January 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300381900321.

Rao:1980:RHH

[RJY80] K. R. Rao, A. Jalali, and P. Yip. Rationalized Hadamard– Haar transform. Applied Mathematics and Computation, 6(3): 263–281, ???? 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900557.

Rubbert:1982:PCS

[RL82] P. E. Rubbert and K. D. Lee.

Patched coordinate systems. Applied Mathematics and Computation, 10–11:235–252, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901928.

Roache:1982:SMS

[Roa82] Patrick J. Roache. Semidirect/marching solutions and elliptic grid generation. Applied Mathematics and Computation, 10–11:729–737, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902181.

Roberts:1982:ATG

[Rob82] A. Roberts. Automatic topology generation and generalised B spline mapping. Applied Mathematics and Computation, 10–11:465–486, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038290203X.

Rajagopal:1984:NSP

[RUS^{+}84] A. K. Rajagopal, V. R. R. Uppuluri, David S. Scott, S. Sitharama Iyengar, and Mohan Yellayi. New structural properties of strings generated by leading digits of 2 ^{N}. Applied Mathematics and Computation, 14(3):221–244, April 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900237.

Sasagawa:1983:ACN

[Sas83] T. Sasagawa. Analytical construction of a negative semidefinite solution of algebraic Riccati equations and its application. Applied Mathematics and Computation, 13(1–2):37–54, August 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300383900292.

Saupe:1983:GBP

[Sau83] Dietmar Saupe. Global bifurcation of periodic solutions to some autonomous differential delay equations. Applied Mathematics and Computation, 13(1– 2):185–211, August 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300383900371.

Saltzman:1982:AGV

[SB82] Jeffrey Saltzman and Jeremiah Brackbill. Applications and generalizations of variational methods for generating adaptive meshes. Applied Mathematics and Computation, 10–11:865– 884, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0096300382902272.

Shieh:1982:TDG

[Shi82] C. F. Shieh. Three-dimensional grid generation using Poisson equations. Applied Mathematics and Computation, 10–11:687– 694, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0096300382902156.

Samimi:1983:GUC

[SL83] Mansour Samimi and V. Lakshmikantham. General uniqueness criteria for ordinary differential equations. Applied Mathematics and Computation, 12 (1):77–88, February 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900437.

Smith:1982:AGG

[Smi82]  Robert E. Smith.  Alge-

braic grid generation. Applied Mathematics and Computation, 10–11:137–170, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901904.

Smith:1984:CPC

[Smi84a] Tony E. Smith. A choice probability characterization of generalized extreme value models. Applied Mathematics and Computation, 14(1):35–62, January 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900456.

Smith:1984:SCC

[Smi84b] Tony E. Smith. A solution condition for complementarity problems: with an application to spatial price equilibrium. Applied Mathematics and Computation, 15(1):61–69, July 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900535.

Sorenson:1982:GGE

[Sor82] Reese L. Sorenson. Grid generation by elliptic partial differential equations for a tri-element Augmentor-Wing airfoil. Applied Mathematics and Computation, 10–11:653–665, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0096300382902132.

Shubin:1982:TDG

[SSB82] G. R. Shubin, A. B. Stephens, and J. B. Bell. Three dimensional grid generation using biharmonics. Applied Mathematics and Computation, 10–11:761–774, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038290220X.

Steger:1982:ABC

[Ste82] Joseph L. Steger. On application of body conforming curvilinear grids for finite difference solution of external flow. Applied Mathematics and Computation, 10–11:295–316, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382901953.

Stewart:1984:BEC

[Ste84] Ian Stewart. Beyond elementary catastrophe theory. Applied Mathematics and Computation, 14(1):25–31, January 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900432.

Stuben:1983:AMA

[Stü83] Klaus Stüben. Algebraic multigrid (AMG): experiences and comparisons. Applied Mathematics and Computation, 13(3– 4):419–451, ???? 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900231.

Sugisaka:1983:DLL

[Sug83] Masanori Sugisaka. The design of on-line linear leastsquares estimators given covariance specifications via an imbedding method. Applied Mathematics and Computation, 13(1– 2):55–85, August 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900309.

Sumita:1984:MLT

[Sum84]  Ushio Sumita.  The matrix

Laguerre transform. Applied Mathematics and Computation, 15(1):1–28, July 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038490050X.

Shendge:1983:CRF

[SV83] G. R. Shendge and A. S. Vatsala. Comparison results for first and second order boundary value problems at resonance. Applied Mathematics and Computation, 12(4):367–380, July 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900474.

Shendge:1984:CRP

[SV84] G. R. Shendge and A. S. Vatsala. Comparison results for parabolic differential equations at resonance. Applied Mathematics and Computation, 15(3): 209–219, October 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038490002X.

Tesfatsion:1980:MSR

[Tes80] Leigh Tesfatsion. C^{3} modeling with symmetrical rationality. Applied Mathematics and Computation, 6(1):51–61, January 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300380900156.

Thames:1982:GTD

[Tha82]  Frank C. Thames.  Gen-

eration of three-dimensional boundary-fitted curvilinear coordinate systems for wing/wingtip geometries using the elliptic solver method. Applied Mathematics and Computation, 10–11:695–716, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902168.

Thomas:1982:NGC

[Tho82a] P. D. Thomas. Numerical generation of composite three dimensional grids by quasilinear elliptic systems. Applied Mathematics and Computation, 10–11:667–686, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382902144.

Thompson:1982:EGG

[Tho82b] Joe F. Thompson. Elliptic grid generation. Applied Mathematics and Computation, 10–11:79–105, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300382901886.

Thompson:1982:F

[Tho82c] Joe F. Thompson. Foreword. Applied Mathematics and Computation, 10–11:xix–xxv, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382901849.

Thompson:1982:GCC

[Tho82d] Joe F. Thompson. General curvilinear coordinate systems. Applied Mathematics and Computation, 10–11:1–30, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300382901850.

Troesch:1980:LVS

[Tro80] B. A. Troesch. The limiting vortex in the similarity solution for a swirling flow. Applied Mathematics and Computation, 6(2):133–144, March 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900387.

Theocaris:1984:NSS

[TT84] P. S. Theocaris and G. Tsamasphyros. A numerical solution of singular integrodifferential equations with variable coefficients. Applied Mathematics and Computation, 15(1):47–59, July 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300384900523.

Tishler:1983:MMA

[TZ83] A. Tishler and I. Zang. A minmax algorithm for non-linear regression models. Applied Mathematics and Computation, 13(1– 2):95–115, August 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900322.

Ueno:1980:IST

[Uen80] Sueo Ueno. Inversion of solar transmitted radiance profiles for the atmospheric optical thickness. Applied Mathematics and Computation, 7(2):

177–186, September 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900041.

Ueno:1983:IVS

[Uen83] Sueo Ueno. An initial-value solution of the least-squares estimation problem with degenerate covariance. Applied Mathematics and Computation, 13(1– 2):87–94, August 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900310.

Ulwadia:1981:NRR

[Ulw81] F. E. Ulwadia. A note on the reliability of redundant systems. Applied Mathematics and Computation, 8(1):51–78, January 1981. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300381900345.

VanRosendale:1983:ADS

[Van83] John Van Rosendale. Algorithms and data structures for adaptive multigrid elliptic solvers. Applied Mathematics and Computation, 13(3–4):453–470, ???? 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900243.

Vargas:1980:NEC

[Var80] Luis G. Vargas. A note on the eigenvalue consistency index. Applied Mathematics and Computation, 7(3):195–203, October 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300380900430.

Vargas:1983:ASR

[Var83] Luis G. Vargas. Analysis of sensitivity of reciprocal matrices. Applied Mathematics and Computation, 12(4):301–320, July 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900449.

Vaughn:1980:SLS

[Vau80] R. L. Vaughn. Solutions of linear Sobolev differential equations. Applied Mathematics and Computation, 6(3):211–216, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/009630038090051X.

Visentin:1980:LPS

[Vis80] F. Visentin. Limiting processes in the stability problem. Applied Mathematics and Computation, 7(1):81–91, July 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300380900351.

Vincent:1983:POD

[VL83] T. L. Vincent and C. S. Lee. Parametric optimization with dynamic systems. Applied Mathematics and Computation, 12 (2–3):169–185, May 1983. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/009630038390005X.

Vatsala:1980:ECR

[VV80] A. S. Vatsala and R. L. Vaughn. Existence and comparison results for differential equations of Sobolev type. Applied Mathematics and Computation, 6(2): 177–187, March 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300380900417.

Wang:1980:SAI

[Wan80] A. P. Wang. On surface area integration and the related mapping. Applied Mathematics and Computation, 6(3):251–262, ???? 1980. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300380900545.

Wang:1981:HARa

[Wan81a] Zhong-De Wang. Harmonic analysis with a real frequency function — I. Aperiodic case. Applied Mathematics and Computation, 9(1):53–73, July 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381901259.

Wang:1981:HARb

[Wan81b] Zhong-De Wang. Harmonic analysis with a real function of frequency — II. periodic and bounded cases. Applied Mathematics and Computation, 9(3): 153–163, October 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900266.

Warsi:1982:BDM

[War82] Z. U. A. Warsi. Basic differential models for coordinate generation. Applied Mathematics and Computation, 10–11:41–77, 1982. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0096300382901874.

Watson:1981:EAC

[Wat81] Layne T. Watson. Engineering applications of the Chow– Yorke algorithm. Applied Mathematics and Computation, 9(2): 111–133, September 1981. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300381900102.

Weibull:1984:SFA

[Wei84] Jörgen W. Weibull. A stockflow approach to general equilibrium with trade frictions. Applied Mathematics and Computation, 14(1):63–76, January 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900468.

Wilcox:1984:MGC

[WH84] Ralph M. Wilcox and Leo P. Harten. MACSYMA-generated closed-form solutions to some matrix Riccati equations. Applied Mathematics and Computation, 14(2):149–166, February 1984. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0096300384900353.

Wong:1980:EAP

[WI80] K. C. Wong and S. Sitharama Iyengar. An efficient algorithm for product computations on computer. Applied Mathematics and Computation, 6 (1):1–5, January 1980. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300380900119.

Watts:1983:SCV

[WSL83] H. A. Watts, M. R. Scott, and M. E. Lord. Solving complexvalued differential systems. Applied Mathematics and Computation, 12(4):381–394, July 1983. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300383900486.

Warsi:1982:NGT

[WZ82] Z. U. A. Warsi and J. P.

Ziebarth. Numerical generation of three-dimensional coordinates between bodies of arbitrary shapes. Applied Mathematics and Computation, 10–11:717–728, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/009630038290217X.

Yagla:1982:TPC

[Yag82] Jon J. Yagla. Test problems, coordinate transformations, and technique for nonsteady compressible flow analysis. Applied Mathematics and Computation, 10–11:525–546, 1982. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0096300382902065.

Yakowitz:1984:CAD

[YR84] Sidney Yakowitz and Brian Rutherford. Computational aspects of discrete-time optimal control. Applied Mathematics and Computation, 15 (1):29–45, July 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900511.

Zamani:1984:CFE

[Zam84] Nader G. Zamani. A collocation finite element method for integration of the boundary layer equations. Applied Mathematics and Computation, 15 (2):109–122, August 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900110.

Zielinski:1984:AFP

[Zie84] Przemyslaw Zielinski. An application of the Fokker–Planck equation in stochastic reservoir theory. Applied Mathematics and Computation, 15(2): 123–136, August 1984. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0096300384900122.